OpenStudy (just_one_last_goodbye):

is the simplified form of 2√3 - 2√3 rational?

OpenStudy (zzr0ck3r):

a-a = 0 is o rational?

OpenStudy (anonymous):

a-a=0 medal pzz ;))

OpenStudy (just_one_last_goodbye):

No

OpenStudy (anonymous):

y??

OpenStudy (zzr0ck3r):

sure it is

OpenStudy (zzr0ck3r):

0=0/4

OpenStudy (anonymous):

yes it is

OpenStudy (just_one_last_goodbye):

.o. sorry i see it now

OpenStudy (anonymous):

a-a=0

OpenStudy (anonymous):

see madal plzz

OpenStudy (zzr0ck3r):

a rational number is a number that CAN be written as the ratio of two integers

OpenStudy (welshfella):

yes 0 is rational

OpenStudy (just_one_last_goodbye):

is the simplified form of 2√3 x 2√3 rational?

OpenStudy (zzr0ck3r):

\(2\sqrt{3}*2\sqrt{3} = 4\sqrt{3*3} = 4*3 = 12\)

OpenStudy (zzr0ck3r):

is this rational?

OpenStudy (anonymous):

yes it is

OpenStudy (zzr0ck3r):

\(\checkmark\)

OpenStudy (just_one_last_goodbye):

Thanks so much